A particle of mass 100 gm and charge 2 microcoulomb is released from rest at a distance of 50 cm from a fixed charge of 5 microcoulomb. Find the speed of the particle when its distance from the fixed charge become 3m?
1
Expert's answer
2020-08-05T16:41:51-0400

"W=q_0(\phi_1-\phi_2)=q_0(k\frac{q}{r_1}-k\frac{q}{r_2})=kqq_0(\frac{1}{r_1}-\frac{1}{r_2})"


"W=mv^2/2\to v=\sqrt{2W/m}"


"v=\sqrt{2kqq_0(\frac{1}{r_1}-\frac{1}{r_2})/m}="


"=\sqrt{2\cdot 9\cdot 10^9\cdot 5\cdot10^{-6} \cdot 2\cdot10^{-6}(\frac{1}{0.5}-\frac{1}{3})/0.1}=1.73(m/s)"